Shwonson Industries reported a deferred tax asset of $8.50 million for the year ended December 31, 2020, related to a temporary difference of $34 million. The tax rate was 25%. The temporary difference is expected to reverse in 2022, at which time the deferred tax asset will reduce taxable income. There are no other temporary differences in 2020–2022. Assume a new tax law is enacted in 2021 that causes the tax rate to change from 25% to 15% beginning in 2022. (The rate remains 25% for 2021 taxes.) Taxable income in 2021 is $44 million.

Required:

1. Prepare the appropriate journal entry to record Shwonson’s income tax expense in 2021.

2. What effect, will enacting the change in the 2022 tax rate, have on Shwonson’s 2021 net income?
Answer #1

Solution:

Requirement 1:

Journal Entry to record Shwonson’s income tax expense in 2021.

Transaction General Journal Debit Credit
1 Income Tax Expense [$ 3.4 M + $ 11 M] $             14.40
Deferred Tax Asset [Note: 1] $             3.40
Income Tax Payable $          11.00
[ To record income tax expense]

Notes:

1) Deferred Tax Asset reversal = $ 34 M * ( 25% - 15%) = $ 3.4 Millions.

2) Income Tax Payable = Taxable Income in 2021 * 25% =  $ 44 M * 25% = $ 11 Millions

Requirement 2:

1) Decrease in Net income = $ 3.4 Millions

2)With the change in tax rate [15% - 25%], there will be decrease in net income due to increase in tax expense of $ 3.4 Millions.

Explanation:

Change in tax expense = $ 34 M * ( 25% - 15%) = $ 3.4 Millions.